# captionUnknown document class (or package)

### v1

@@ 4.2 LLMs for Enhancing 3D Task Performance
3D Component LLM Component 3D+LLM Method Subsection 3D Geometry Vision Model Tuning LLM Abilities Fine-tuning LLM Base # Parameters Hardware Alignment Module Date Chen et al. [274] 3.2 SG CLIP F WK None GPT2 1.5B 1 RTX3080 - 09/22 ConceptFusion [18] 3.2/3.4 RGB-D OpenSeg F IF/R None GPT3 175B 1 3090 - 02/23 ViewRefer [20] 3.2 MVI Multi-View Transformer T WK None GPT3 175B 4 A100 Transformer 03/23 LLM-Grounder [21] 3.2 PC/NeRF OpenScene/LERF F IF/ICL/R None GPT3.5/4 - - - 09/23 Abdelreheem et al. [22] 3.2 Mesh Vision Transformer F IF/WK None GPT3.5 - 1 3090 - 09/23 Transcribe3D [23] 3.2 PC Groupfree Transformer F IF/R None GPT3.5/4 - - - 10/23 zero-shot 3DVG [16] 3.2 RGB-D Mask3d F IF/ICL/R None GPT3.5/4 - - Transformer 11/23 3DAP [273] 3.2 MVI - - IF/ICL PT GPT4V - - - 12/23 3D-LLM [153] 3.3 PC Mask2Former/SAM F IF/ICL/R LF OPT/Flan-T5 9B/2.7B/3B 64 V100 QFormer 07/23 Chat-3D [171] 3.3 PC Point-BERT F IF/ICL/R AF Vicuna 7B - Linear layer 08/23 LEO [270] 3.3 PC/MVI OpenClip/PointNet++ F IF/ICL/R LoRA Vicuna 7B 8 A100 Transformer 11/23 LL3DA [19] 3.3 PC/3D-BB ScanNet scene encoder F IF/ICL/R AF OPT 1.3B 8 3090 QFormer 11/23 Point-LLM [267] 3.3 PC Point-BERT F IF/ICL/R Full LLaMA 7B/13B 8 A100 Linear layer 12/23 GPT4Point [268] 3.3 PC Point-BERT F IF/ICL/R LF OPT/Flan-T5 6.7B/3B 8 A100 QFormer 12/23 Chat-3D v2 [172] 3.3 PC Uni-3D F IF/ICL/R AF Vicuna 7B 4 A40 MLP 12/23 LiDAR-LLM [271] 3.3 PC/VM VoxelNet F IF/ICL/R AF LLaMA 7B 4 A100 Transformer 12/23 3DMIT [269] 3.3 PC EPCL/Uni3D F IF/ICL/R LoRA Vicuna 7B 8 A100 Linear layer 01/24 Scene-LLM [266] 3.3 PC/VM ConceptFusion F IF/ICL/R LF LLaMA-2 7B 32 A100 Linear layer 03/24 Point-Bind [272] 3.4 PC I2P-MAE F IF/ICL/R LF LLaMA 7B 8 A100 Linear layer 09/23 JM3D-LLM [279] 3.4 PC PointNet++/PointMLP/Point-BERT F IF/ICL/R LF Vicuna 7B 3 A100 MLP 10/23 LLM-Planner [12] 3.5 VM Seg&Depth F IF/ICL/R None GPT3 175B - - 03/23 SayPlan [275] 3.5 SG SG generator F IF/ICL/R None GPT4 - - - 07/23 VoxPoser [13] 3.5 RGB-D OWL-ViT/SAM F IF/ICL/R None GPT4 - - - 07/23 UniHSI [277] 3.5 RGB-D - - IF/ICL/R None GPT3.5/4 - 1 A100 - 09/23 LAN-grasp [14] 3.5 MVI OWL-ViT F IF/ICL/R None GPT4 - - - 10/23 Agent3D-Zero [17] 3.3/3.5 MVI - - IF/ICL/R PT GPT4V - - - 03/24 NaviLLM [11] 3.5 MVI EVA-CLIP-Large F IF/ICL/R LF Vicuna 7B 8 A100 Transformer 12/23 MultiPLY [24] 3.3/3.4/3.5 PC ConceptGraph T IF/ICL/R AF Vicuna 13B 128 V100 Linear layer 01/24 ManipLLM [15] 3.5 RGB-D CLIP F IF/ICL/R AF LLaMA 7B 1 A100 Linear layer 03/24 3D-VLA [276] 3.3/3.5 PC/MVI Mask2Former/SAM F IF/ICL/R LoRA Flan-T5 3B 384 V100 QFormer 03/24 PolyGen [280] 3.6 n-gon Mesh Custom Transformer F - Full Custom - 4 V100 - 02/20 LLMR [281] 3.6 Unity, Mesh Dall-E-2, CLIP F IF/ICL/R None GPT4 - 1 3080 - 09/23 3D-GPT [26] 3.6 Blender, Mesh - - IF/R None GPT3.5/4 - - - 10/23 MeshGPT [282] 3.6 Mesh GPT-2-style Transformer F - Full GPT-2-style $\sim$345M 4 A100 - 11/23 ShapeGPT [283] 3.6 SDF 3D VQ-VAE, CLIP F IF/R Full T5 - 4 A100 T5 11/23 Holodeck [284] 3.6 Mesh - - IF/R None GPT4 - 8 RTX 8k CLIP/SBERT 12/23 GALA-3D [25] 3.6 3DGS MVDream, Control-Net F IF/R None Any (GPT3.5) - - - 02/24
TABLE I: Summarization of 3D-LLMs methods. The 3D Geometry column collects the 3D geometric information being used for each method, such as Point Cloud (PC), Multi-view Images (MVI), RGB+Depth (RGB-D), 3D Bounding Box (3D-BB), Scene Graph (SG), Voxel Map (VM) and NeRF. The tuning column outlines whether the vision model is finetuned during training (True/False). IF, ICL, R, and WK denote the LLM Abilities of Instruction-following, In-Context Learning, Reasoning, and World Knowledge. The Fine-tuning column summarizes how LLM components are fine-tuned, such as Prompt Tuning (PT), Low-Rank Adaptation (LoRA), Adaptive Fine-tuning (AF), Layer Freezing (LF), or Full Fine-tuning (Full). The Hardware column shows methods that are trained either with numbers of Nvidia GPU and the specific GPU type or with no training involved.
@@ 4.2.1 Knowledge-enhanced approaches
There are several methods that utilize LLM world knowledge. Chen et al. [274] use LLMs for 3D room classification from RGB-D images. Here, the knowledge embedded in LLMs is used to determine the room category based on the object category information contained in the room. Firstly, this approach creates scene graphs from Matterport3D [285] data with nodes for regions and objects, and object nodes linked to room nodes. Next, key objects are selected to form queries per room type. LLMs score descriptions extracted from the selected objects, with the highest score predicting the room label. Spatial information like sizes or locations can also be provided.
@@
ViewRefer [20] uses LLMs to expand the grounding texts with view-related descriptions. For instance, given the original text “Facing the front of the couch, the table that is to the right of the couch”, the LLM is used to create a similar sentence but from another speaker’s point of view, e.g. “With back to the front of the couch, pick the table that is on the left side of the couch”. By rephrasing input texts and their opposite-view synonyms several times, the model improves cross-view grounding. It also adopts a fusion transformer with inter-view attention and includes learnable multi-view prototypes, which capture inter-view knowledge and further boost 3D grounding performance.
Abdelreheem et al. [22] address the problem of semantic correspondence in 3D shapes. They classify 3D shapes by feeding rendered views into a BLIP2 model to generate lists of class proposals. ChatGPT [286] unifies these into a single class per shape. ChatGPT also generates semantic part names and pairwise mappings (e.g. arm $\rightarrow$ wing). A 3D segmenter then segments shapes based on semantic regions, utilizing the part mappings to produce sparse correspondence maps.
@@
The knowledge-enhanced strategies described above enable strong performance especially in zero-shot scenarios where no labeled 3D data is available for a particular object or scene type. This allows open-ended reasoning about object parts, relationships, and semantics beyond fixed ontologies, as demonstrated by (i) Chen et al. [274] generating spatial and semantic object descriptions, (ii) ViewRefer  [20] describing multi-view object relationships, and (iii) Abdelreheem et al. [22] generating and matching object part semantics across shapes.
Method
@@
Captioning
Grounding QA
Situated
QA
Dialogue Planning Navigation Manipulation
@@
4.2.2 Reasoning-enhanced approaches
Along with world knowledge, the reasoning abilities of LLMs also help tackle other 3D tasks, particularly visual grounding in complicated 3D scenes with detailed geometry and multiple objects. In such cases, text descriptions of the objects should include their appearance and spatial relationship with surrounding items. Ordinary grounding methods [287] often struggle in this setting due to their inability to understand detailed text descriptions. LLM-Grounder [21], Transcribe3D [23], and zero-shot 3DVG [16] approach this problem by leveraging the reasoning capability of LLMs to analyze the text descriptions and generate a sequence of instructions for locating the objects using existing grounding toolboxes. Specifically, an LLM first identifies anchor and target objects from the text descriptions. It then analyzes the spatial relationship (or described attributes) between multiple candidate objects based on their coordinates returned by the grounding tool to select the candidate that best matches the text description. In addition, (i) Transcribe3D [23] and LLM-Grounder [21] employ a multi-round interactive question-and-answer process to help users clarify their intentions, prompting them to provide instructions with more information to arrive at more accurate results, whereas (ii) LLM-Grounder includes multiple grounding tool choices, such as OpenScene [29] or LERF [30] to accommodate different 3D representations like point clouds or NeRF. A common drawback of these approaches is the “blindness” of the LLM since it is provided only with abstract textual descriptions of the 3D scene instead of the original point cloud of the scene. This may lead to the loss of crucial scene details. Consequently, when a 3D scene contains multiple objects of the same class, the absence of necessary scene details means that ambiguities of the text-based references cannot be resolved, which limits overall performance.
Besides visual grounding, the reasoning capability of LLMs facilitates other tasks as well. 3DAP [273] leverages GPT-4V to infer 3D information of an object from its 2D images using a visual prompting technique, where it annotates the input image with a 3D axis to enhance the LLM’s awareness of the 3D scale. ConceptFusion [18] uses GPT3 to generate instructions using pre-defined elementary spatial comparison modules to enable more complex spatial reasoning using their proposed 3D feature map.
4.3 LLMs for 3D Multi-Task Learning
Many works focus on using the instruction-following and in-context learning capabilities of LLMs to unify multiple 3D tasks into a single language space. By employing different text prompts to denote different tasks, these studies aim for LLMs to serve as a unifying dialogue interface. Achieving multi-task learning using an LLM typically involves several key steps, beginning with the construction of 3D-text data pairs [19, 153, 270]. These pairs require crafting task instructions in text form and defining the output for each different task. Next, 3D data (typically in the form of point clouds) are fed into 3D encoders [288, 34] to extract 3D features. Alignment modules [171, 153, 172, 266] are subsequently used to (i) align 3D features with text embeddings from the LLMs across multiple levels (object-level, relationship-level, and scene-level) and (ii) translate the 3D features into tokens interpretable by the LLMs. Finally, an appropriate training strategy [272, 19, 153, 270, 269] needs to be selected, such as single or multi-stage 3D-language alignment training and multi-task instruction fine-tuning.
In the remainder of this section, we explore these aspects in detail. We additionally summarize the scope and capabilities of each method reviewed in this section in Table II.
@@ 4.3.1 Data for Multi-Task Learning
As demonstrated in Table II, we divide tasks into four categories: captioning, grounding, question answering (QA), and embodied agent tasks (i.e., planning, navigation, and manipulation). Accordingly, each task’s textual output follows a predefined format. For captioning and QA tasks, the output is plain text and is not constrained by a specific format. The output of the grounding task is a 3D bounding box, typically the coordinates of the referred object’s center along with its 3D size. Normally, the values for point and size are normalized to fall within the range of 0-255 [19] which limits the range of tokens the LLM is required to predict. For planning, the model outputs a series of steps to execute a task in text form, whereas for navigation, the output is a series of spatial coordinates. For manipulation, the output is action sequences in text form. Existing methods follow these guidelines to construct their multi-task instruction fine-tuning datasets.
@@
Once the text format has been decided upon, different methods use different strategies to obtain text annotations for their datasets. Several approaches utilize human labelers to generate ‘ground truth’ annotations for each sample [218, 228, 234, 233], however this can be an expensive and time-consuming process. Another approach is to use ChatGPT [286] to generate text annotations for each sample, which is a strategy utilized by 3DMIT [269], LiDAR-LLM [271], Chat-3D [171] and Chat-3D v2 [172]. Here, 3D scene data is converted into text (usually by describing object bounding boxes and spatial relations textually), and a task description is created to describe the desired output. To guide ChatGPT towards the expected output format for the task, demonstration examples are provided, which allows ChatGPT to perform in-context learning to generate plausible text annotations for other 3D scenes. Alternatively, other multi-task datasets [19, 266] are constructed simply be merging existing 3D Vision-Language (VL) datasets [218, 228, 234, 153]. Some multi-task datasets are constructed using a combination of these three approaches, as seen in LEO [270], LiDAR-LLM [271] and 3D-LLM [153], seeking to combine the accuracy of human annotations with the scalability provided by using LLM-generated annotations.
4.3.2 Training an LLM for multiple 3D tasks
The first step in training an LLM for multiple 3D tasks involves obtaining meaningful 3D features, where the extraction method varies depending on the type of 3D scene. For individual object point clouds, Point-LLM [267], Chat-3D [171], and GPT4Point [268] utilize Point-BERT [288] to extract 3D object features. For indoor scenes, LEO [270] uses PointNet++ [34] for feature extraction, whereas Chat-3D v2 [172] and 3DMIT [269] segment the scene and use Uni-3D [289] to extract features for each segmented part. Meanwhile, MultiPLY [24] integrates the extracted object features into a scene graph [290] to represent the entire scene. 3D-LLM [153] and Scene-LLM [266] lift features from 2D multi-view images into a 3D representation. 3D-LLM [153] lifts 2D semantic features from Mask2Former [291] or SAM [211]. Scene-LLM [266] follows ConceptFusion [18] to fuse both global information and local details, mapping pixel-wise CLIP features into point-wise 3D features. For outdoor 3D scenes, LiDAR-LLM [271] uses VoxelNet [292] to extract 3D voxel features.
For alignment modules, as discussed in Sec. 4.1, various network architectures are used. Notably, MultiPLY [24] employs distinct linear layers to align features from each modality. Chat-3D [171] and Chat-3D v2 [172] use a one-layer vanilla transformer to allow 3D object features to attend to each other during alignment. LEO [270] and LiDAR-LLM [271] utilize modified transformers as their alignment modules to better suit different types of 3D data (dense point clouds vs. sparse LiDAR). LEO [270] modifies the self-attention mechanism to explicitly encode spatial relations between object pairs in point clouds. In contrast, LiDAR-LLM [271] employs both self-attention and cross-attention mechanisms to align Bird’s Eye View (BEV) features with text features. 3D-LLM [153] and GPT4Point [268] adopt Q-Former, while LL3DA[19] adds an additional branch on top of Q-Former that allows query tokens to interact with visual prompts provided by users.

### v2

@@ 4.2.1 Knowledge-enhanced approaches
There are several methods that utilize LLM world knowledge. Chen et al. [148] use LLMs for 3D room classification from RGB-D images. Here, the knowledge embedded in LLMs is used to determine the room category based on the object category information contained in the room. Firstly, this approach creates scene graphs from Matterport3D [190] data with nodes for regions and objects, with object nodes linked to room nodes. Next, key objects are selected to form queries per room type. LLMs score descriptions extracted from the selected objects, with the highest score predicting the room label.
@@
ViewRefer [20] uses LLMs to expand the grounding texts with view-related descriptions. For instance, given the original text “Facing the front of the couch, the table that is to the right of the couch”, the LLM is used to create a similar sentence but from another speaker’s point of view, e.g. “With back to the front of the couch, pick the table that is on the left side of the couch”. By rephrasing input texts and their opposite-view synonyms several times, the model improves cross-view grounding. Abdelreheem et al. [22] address the problem of semantic correspondence in 3D shapes by using a BLIP2 model to generate lists of text class proposals from different rendered views of a 3D shape. ChatGPT [191] then unifies these proposals into a single class proposal for the entire object.
@@
The knowledge-enhanced strategies described above enable strong performance, especially in zero-shot scenarios where no labeled 3D data is available for a particular object or scene type. This allows open-ended reasoning about object parts, relationships, and semantics beyond fixed ontologies, as demonstrated by (i) Chen et al. [148] generating spatial and semantic object descriptions, (ii) ViewRefer [20] describing multi-view object relationships, and (iii) Abdelreheem et al. [22] generating and matching object part semantics across shapes.
4.2.2 Reasoning-enhanced approaches
Along with world knowledge, the reasoning abilities of LLMs also help tackle other 3D tasks, particularly visual grounding in complicated 3D scenes with detailed geometry and multiple objects. In such cases, text descriptions of the objects should include their appearance and spatial relationship with surrounding items. Ordinary grounding methods [192] often struggle in this setting due to their inability to understand detailed text descriptions. LLM-Grounder [21], Transcribe3D [23], and zero-shot 3DVG [16] approach this problem by leveraging the reasoning capability of LLMs to analyze the text descriptions and generate a sequence of instructions for locating the objects using existing grounding toolboxes. Specifically, an LLM first identifies anchor and target objects from the text descriptions. It then analyzes the spatial relationship (or described attributes) between multiple candidate objects based on their coordinates returned by the grounding tool to select the candidate that best matches the text description.
A common drawback of these approaches is the “blindness” of the LLM, since it is provided only with abstract textual descriptions of the 3D scene instead of the original point cloud. This may lead to the loss of crucial scene details. Consequently, when a 3D scene contains multiple objects of the same class, the absence of necessary scene details means that ambiguities of the text-based references cannot be resolved, which limits overall performance.
Besides visual grounding, the reasoning capability of LLMs can facilitate other tasks. 3DAP [149] leverages GPT-4V to infer 3D information of an object from its 2D images using a visual prompting technique, where it annotates the input image with a 3D axis to enhance the LLM’s awareness of the 3D scale. ConceptFusion [18] uses GPT3 to generate instructions using pre-defined elementary spatial comparison modules to enable more complex spatial reasoning using their proposed 3D feature map. SpatialPIN [150] utilizes GPT-4V to infer interactive objects, potential tasks and plan of actions from a single RGB image.
4.3 LLMs for 3D Multi-Task Learning
Method 3D Captioning 3D Grounding 3D Conversation 3D Embodied Agents 3D Generation
@@ 4.3 LLMs for 3D Multi-Task Learning
Single
Multi
QA Situated-QA Dialogue Plan. Nav. Manip. object
@@
Spartun3D-LLM [193]
Chat-Scene [178]
LLaVA-3D [162]
Robin-3D [163]
Video-3D LLM [165]
GPT4Scene [166]
LSceneLLM [167]
Ross3D [170]
3D-LLaVA [171]
VG-LLM [173]
3DRS [176]
@@
Many works focus on using the instruction-following and in-context learning capabilities of LLMs to unify multiple 3D tasks into a single language space. By using different text prompts to denote different tasks, these works allow for LLMs to serve as a unifying dialogue interface. Achieving multi-task learning using an LLM typically involves several key steps, beginning with the construction of 3D-text data pairs [19, 115, 152]. These pairs require crafting task instructions in text form and defining the output for each different task. Next, 3D data (typically in the form of point clouds) are fed into 3D encoders [194, 195] to extract 3D features. Alignment modules [80, 115, 81, 147] are subsequently used to (i) align 3D features with text embeddings from the LLMs across multiple levels (object-level, relationship-level, and scene-level) and (ii) translate the 3D features into tokens interpretable by the LLMs. Finally, an appropriate training strategy [19, 115, 177, 152, 157, 196] needs to be selected, such as single or multi-stage 3D-language alignment training and multi-task instruction fine-tuning. In the remainder of this section, we explore these aspects in detail. We additionally summarize the scope and capabilities of each method reviewed in this section in Table II.
@@ 4.3.1 Data for Multi-Task Learning
As shown in Fig. 2, we divide tasks into five categories: captioning, grounding, question answering (QA), embodied agent tasks (i.e., planning, navigation, and manipulation), and generation. Accordingly, each task’s textual output follows a predefined format. For captioning and QA tasks, the output is plain text and is not constrained by a specific format. The output of the grounding task is a 3D bounding box, typically the coordinates of the referred object’s center along with its 3D size. Normally, the values for point and size are normalized to fall within the range of 0-255 [19] which limits the range of tokens the LLM is required to predict. For planning, the model outputs a series of steps to execute a task in text form, whereas for navigation, the output is a series of spatial coordinates. For manipulation, the output is action sequences in text form. Existing methods follow these guidelines to construct their multi-task instruction fine-tuning datasets.
@@
Once the text format has been decided upon, different works use different strategies to obtain text annotations for their datasets, constructing 3D-text pairs for training. Several approaches ask human labelers to generate ‘ground truth’ annotations for each sample [101, 111, 58, 59]. However this can be an expensive and time-consuming process. Another approach [157, 155, 80, 158, 81] is to use an LLM (e.g. ChatGPT [191]) to generate text annotations for each sample. Here, 3D scene data is converted into text (usually by describing object bounding boxes and spatial relations textually), and a task description is created to describe the desired output. To guide ChatGPT towards the expected output format for the task, demonstration examples are provided, which allows ChatGPT to perform in-context learning to generate plausible text annotations for other 3D scenes. Alternatively, other multi-task datasets [19, 147] are constructed by merging existing 3D Vision-Language (VL) datasets [101, 111, 58, 115]. Several recent multi-task datasets adopt a hybrid strategy that combines these three approaches [152, 155, 115], aiming to leverage both the accuracy of human annotations and the scalability of LLM-generated ones.
4.3.2 3D Feature Tokenization for LLM training
The first step in training an LLM for multi-task 3D understanding lies in extracting semantically and spatially meaningful 3D features. We group existing methods into two main categories depending on their input modality and spatial encoding strategies.
Point-based methods operate directly on point clouds, offering explicit spatial grounding and high geometric precision. They can be further divided into: (1) Scene-centric models [115, 162, 19, 154, 157, 158, 153], which process raw scene point clouds using pretrained or custom point encoders such as Point-BERT [194] to obtain the 3D scene features, then alignment module (such as MLP) is used to project 3D features to 3D tokens; (2) Object-centric models [152, 80, 81, 156, 168, 163, 158, 153, 159, 161], instead of processing the whole 3D scenes, each 3D object is segmented via 3D segmentation, such as Mask3D [197] and then encoded to 3D object feature. Additional 3D positional embeddings are incorporated into each object feature to generate the final 3D tokens. While effective at modeling discrete entities, they often lack comprehensive scene-level spatial understanding; and (3) Voxel-centric methods [147, 155], which discretize the 3D space into voxel grids and apply an encoder such as VoxelNet [198] to get 3d voxel feature. Despite the advantage of maintaining spatial structures and metric accuracy, these methods generally suffer from limited semantic richness due to annotation scarcity and the sparsity of 3D data. Moreover, due to the relatively small scale of existing point cloud datasets, 3D-LLMs trained solely on point-based representations typically underperform their 2D counterparts in terms of general-purpose semantic understanding [199].
Multi-view image-based methods directly encode multi-view images of a 3D scene using a 2D image encoder, and then lift the 2D image features into 3D space to obtain 3D tokens. The advantage of this paradigm is that it combines the semantic richness and well-aligned text-image representation of large-scale 2D vision-language models (VLMs) with 3D spatial reasoning. These can be categorized as: (1) Projection-based methods such as LEO-VL [175], which project 2D pixel features from 2D images into 3D coordinates using calibrated camera intrinsics, thus obtaining a feature for each 3D point. By further embedding the XYZ coordinates through 3D positional encoding, the final 3D tokens are generated; (2) Reconstruction-based methods [169, 160], which first reconstruct a 3D scene (e.g., via 3D gaussian splatting or NeRF) and then align it with language; and (3) Geometry-augmented methods [173, 172, 174], which use VGGT [200]-like encoder to generate both geometry-aware and semantic representations without explicit projection or reconstruction. These approaches benefit from pretrained VLMs and rich 2D image-text datasets, yet the additional lifting step can introduce projection errors and lead to imprecise spatial relationships in the constructed 3D space. (4) Temporal-augmented methods [162, 165, 166, 170, 176] repurpose video streams as implicit 3D observations, leveraging temporal continuity as a proxy for spatial structure. these methods feed video frames to existing 2D-VLMs, without architectural modifications, directly inherit their strong language grounding and reasoning capabilities. Additionally, videos are widely available, and adjacent frames are naturally correlated in both time and space, implicitly encoding 3D geometry through motion. This category currently yields the best empirical performance among 3D-LLMs due to the scalability and semantic richness of video-text data, as shown in Table IV.
4.3.3 Training Strategy for Multi-Task Learning